HUBBLE SPACE TELESCOPE DAILY REPORT #2569 PERIOD COVERED: 0000Z (UTC) 02/28/00 - 0000Z (UTC) 02/29/00 Daily Status Report as of 060/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 8439 (Bias Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no reported problems. 1.2 Completed WF/PC-2 8192 (A Snapshot Survey of Probable Nearby Galaxies) The WF/PC-2 was used to make a snapshot survey images of a probable nearby galaxy KK044. The census of galaxies in the Local Volume of space is very incomplete, particularly at the faint end. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. The proposal completed nominally. 1.4 Completed Three Sets of STIS/CCD 8421 (Spectroscopic Sensitivity for (MAMA and) CCD) The Space Telescope Imaging Spectrograph (CCD) was used to perform basic sensitivity measurement for some supported (MAMA and) CCD spectroscopic modes. Sensitivity measurements are done for a few primary tilts of the gratings. The proposal completed nominally, with no reported problems. 1.5 Completed Two Sets of WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations completed with no reported problems. 1.6 Completed Four Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 8459 (Decontaminations and Associated Observations Pt. 2/2) The WF/PC-2 was used to perform the monthly decons. Also included are instrument monitors tied to decons: photometric stability check, focus monitor, pre- and post-decon internals {bias, intflats, kspots, & darks}, and UV throughput checks. The observations were completed nominally. 1.8 Completed Two Sets of WF/PC-2 8442 (WF/PC-2 Cycle 8 Standard Darks) The WF/PC-2 was used to make a series of Cycle 8 standard dark calibration observations. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2/STIS/CCD 8208 (Galaxy Interactions, Tidal Debris, and the Origin of Intracluster Light in the Coma Cluster) The WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to obtain deep WFPC2 and parallel STIS images of low surface brightness tidal debris that we have recently discovered in the Coma cluster; the material is being stripped from its parent galaxy and added to the general cluster background. These images will enable direct study of the brightest blue and red supergiants, globular clusters, and star forming regions which may be present, or will place strong limits on the numbers of such objects and any recent star formation. The observations completed nominally. 1.10 Completed WF/PC-2 8457 (UV Earth Flats) The WF/PC-2 was used to obtain a series of Ultraviolet Earth flats. The observations were completed with no reported problems. 1.11 Completed WF/PC-2 8445 (WF/PC-2 Cycle 8 Earth Flats) The WF/PC-2 was used to make Cycle 8 Earth flat calibration observations. This is to obtain sequences of earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 7 Successful: 6 The acquisition scheduled for 060/0708Z failed when the search radius limit was exceeded on FGS-3. HSTAR 7525 was written. STIS proposal 8202, which occurs during the period of the next report, was affected. Scheduled Re-acquisitions: 9 Successful: 9 2.2 FHST Updates: Scheduled: 16 Successful: 16 2.3 Operations Notes: Servicing Mission Observatory Verification (SMOV) continues, using CCS string "A" as the prime operational string. Using ROP SR-1A, the SSR EDAC error counter was cleared once. There was an EMC retry at 059/130905Z. ROP NS-12 was utilized to reset the STIS flight software error counter at 059/1340Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of SMOV and normal science observations and calibrations. /CAW